Newhard Company assigns overhead cost to jobs on the basis of 112% of direct labor cost. The job cost sheet for Job 313 includes $20,728 in direct materials cost and $10,600 in direct labor cost. A total of 1,600 units were produced in Job 313 Required: a. What is the total manufacturing cost assigned to Job 313? b. What is the unit product cost for Job 313? a. Total manufacturing cost b. Unit product cost
Answer #1

a) To determine the total manufacturing costs you add the direct materials, the direct labor, and the manufacturing overhead applied. In this case that would be

$20728+$10600+(10600*112%) = $43200

b) Unit product cost = Total manufacturing cost / number of units

= 43200/1600 = $27 per unit

Thank You...